OpenStudy (anonymous):

Find all the zeroes of the equation. x3 + x2 = –3x – 3

OpenStudy (anonymous):

are the first terms 3x+2x??

OpenStudy (anonymous):

no it is x^3+x^2

OpenStudy (anonymous):

first you got set equation = 0 (x^3+x^2)+(3x+3)=0 factor out GCF (x^2+3)(x+1)=0 x^2+3=0, x+1=0 solve for x

OpenStudy (anonymous):

xyou shoudl get x=-1, i\[\sqrt{3}\],-i\[\sqrt{3}\]